Perform an operation on the given system that eliminates the indicated variable. Write the new equivalent system.
−5x + 5y − 4z = 4
10x − 9y + z = −20
−x + 9y + z = 8
Eliminate the x-term from the second equation.

Answer by ikleyn(52781)

.
Nothing can be easier.

Multiply the first equation by 2 and then add the result to the second equation.


After multiplying, you will get 

-10x + 10y - 8z =   8,    (1')
 10x -  9y +  z = -20,    (2')
  -x +  9y +  z =   8.    (3')

Now add (1') and (2') and write the sum as an updated second equation. You will get

-10x + 10y - 8z =   8,    (1'')
  0x +   y - 7z = -12,    (2'')
  -x +  9y +  z =   8.    (3'')

Variable x is just eliminated. You can write your system in the form

-10x + 10y - 8z =   8,    (1'')
         y - 7z = -12,    (2'')
 -x  +  9y +  z =   8.    (3'')

Done and solved.